OpenStudy (senia01):

David has a hose that is 27 ft long. He wants to cut it into two pieces so that the longer piece is 3 ft more than 2 times the length of the shorter piece. How long will each piece of hose be? Shorter piece: ft Longer piece: ft

OpenStudy (anonymous):

Hi there! We need to write a couple of equations so that we can solve this. Let's call the two pieces x and y x + y =27 x = 2y + 3 do you see where i got these equations from?

OpenStudy (senia01):

yes x+y+27 is the shorter piece+longer piece= the hose

OpenStudy (anonymous):

Cool. So now we have to solve these. x = 27 - y substitute this into x = 2y + 3 (27-y)=2y+3

OpenStudy (anonymous):

And solve that equation to find y

OpenStudy (senia01):

8?

OpenStudy (anonymous):

Perfect. Now that you have y, you can find x. x + y = 27 Substitute y = 8 and solve for x

OpenStudy (anonymous):

Let x be the length of the shorter piece of rope. Solve the following for x: x + (2 x + 3) = 27

OpenStudy (senia01):

so the longer piece is 19 and the shorter piece is 8?

OpenStudy (anonymous):

Yes.

OpenStudy (senia01):

Can you help me with one more?

OpenStudy (anonymous):

I'l try.